A wheel rotates clockwise 10 times per second. What is its angular speed? B. What will be its angular displacement after 3 seconds? C. What will be the tangential velocity of the outside of the wheel?
Answer #1

given frequency= 10 Hz

angular frequency= 2*3.14* 10 =62.8 rad/sec.

angular displacement = 62.8 *3= 188.4 rad

tangential velocity = r*ang.freq = 62.8* r m/s.